victoriacollege.instructure.com 20 Question 3 An open top tank is filled with a liquid that has a specific gravity of 1.735. The liq exerts a head pressure of 450 inches w.c. What is the level in inches? 780.7500 Question 4 20 pt An open top tank is filled with a liquid that has a specific gravity of 0.873. The liquid exerts a head pressure of 200 inches w.c. What is the level in inches? 229.0950 Question 5 20 pts A sealed tank has a liquid with a specific gravity of 1.00 in it. The pressure in its vapor space is 1 psi If the height of the liquid is 10.0 feet, how much head pressure in inches w.c., will it exert?
Answer #1

3. open top tank

density of liquid = 1.735*rho ( where rho is density of waqter)

Pressure head = 450 in w.c.

let the level of liquid in inches b e h

p = 450 in wc = 450*rho*g

hence 1.735*rho*g*h = 450*rho*g

h = 259.365 in

4. given

p = 200 in w.c. = 200*rho*g

hence let level of liquid be h

0.873*rho*gh = 200*rho*g

h = 229.095 in

5. specific gravity = 1

Po = 1 psi

h = 10 ft

P = Po + rho*g*h

P = 1 + 0.000036127*386.2205*10*12

P = 2.67435856042 psi

in water column units

P = 0.000036127*386.2205*h = 2.67435856042

h = 191.6692367 in w.c.